The neutralization of a solution of 1.2 g of a substance containing a mixture of H2 C2 O4 .2H2 O, KHC2 O4 .H2 O and different impurities of a neutral salt consumed 18.9 mL of 0.5 N NaOH solution. On titration with KMnO4 solution, 0.4 g, of the same substance needed 21.55 mL of 0.25 N KMnO4. Calculate the % composition of the substance. 
I CHALLENE YOU TO SOLVE IT.. 
IS THERE ANY MAI KA LAL WHO CAN SOLVE IT

Grade:10

1 Answers

Vikas TU
3 years ago
equivalent of h2c2o4.2h2o + equivalent of khc2o4.h2o = equivalent of naoh
let mass of each be x and y
 x/126 + y/146 = 0.00945
146x+126y=173.84
1.2 mixture contains x grams and y grams
0.4 mixture contains x/3 grams and y/3 grams
equivalent of 1 + equivalent of 2 = 0.005387
x/378 + y/ 438 = 0.005387
438x+378y=890.40